OpenStudy (anonymous):

Find the future value of an ordinary annuity that calls for depositing $100 at the end of every 6 months for 5 years into an account that earns 11% interest compounded semiannually.

OpenStudy (anonymous):

please help me thank you :)

OpenStudy (wolf1728):

r = 11/2 = 5.5 Total = {100 * [(1+.055)^11 -1]/.055} -100 1,358.35 I'll work that out just to make sure that really is the correct answer Look at the attachment for the formula.

OpenStudy (wolf1728):

Calculating it for 10 six-month periods + 9 six-month periods + 8 six-month periods, etc totals to 1,369.97 which is slightly higher than 1,358.35 One thing different between the two was that I used 11% compounded semi-annually which = 11.3025 annually.

OpenStudy (wolf1728):

Using 11% compounded semi-annually gives a semi-annual rate of 11.3025/2 = 5.51625. Total = {100 * [(1+.0551625)^11 -1]/.0551625} -100 Total = {805.1480644 / .0551625} -100 = 1,359.59 I recalculated the 10 semi-annual periods of interest and they total 1,359.59 Pretty good huh?

OpenStudy (anonymous):

thank you so much :)